OpenStudy (anonymous):

Find the 29th term of the following sequence. 12, 14, 16, 18, . . .

OpenStudy (anonymous):

it is in arithmetic progression,so 29 th term = a + (29-1) d where a = 12 ,d = 14-12 =2 29 th term =12 + 28 (2) = 68

OpenStudy (anonymous):

Thank you